OpenStudy (anonymous):

What is the slope of the line that contains the points (-1, 8) and (5, -4)? Enter your answer as an integer.

OpenStudy (anonymous):

solpe is \[\frac{ y2-y1 }{ x2-x1 }\] \[\frac{ -4-8 }{ 5-(-1) }\]

OpenStudy (anonymous):

That's right, slope= (y2-y1)/(x2-x1)=(-4-8)/(5-(-1))=-12/6=-2.